# Image Intrinsic Scale Assessment: Bridging the Gap Between Quality and Resolution

### v1

@@ Acknowledgements
In Sec. S2.5, we discuss the precision of subjective measurements, indicated by the confidence intervals of the aggregated annotations. The precision levels for IISA and NR-IQA are comparable. However, the leverage factor $\gamma>1$ implies that minor changes in quality result in larger variations in scale. Therefore, achieving a specific precision in measuring scale equates to an even finer precision in measuring quality. Thus, $\gamma$ acts as a leverage or magnifying factor. This indicates that the sensitivity of our annotation tool for IIS is higher when detecting subtle differences in quality compared to traditional quality rating scales, making IISA suitable for fine-grained quality assessment.
@@
Aiming for higher quality in the Pixabay subset, from a list of 71 camera models, we focused on newer ones released after 2010 with full-frame sensors. From a pool of over one million images on Pixabay with EXIF information, we filtered for photos with a width greater than 4,000 pixels, resulting in approximately 18,000 images that met all criteria. We sampled for diversity 600 images, maintaining uniform distributions among binned normalized favorites, likes, downloads, and user tags using a method similar to [34]. Most of these were captured with Canon EOS 1/5/6D Mark 2/3/4 cameras, using various lenses and capture settings. For the KonIQ-10k subset, we also sampled for diversity regarding quality levels and machine tags – with a confidence greater than 80% – using the same stratified procedure.
The last filtering step we applied was removing images containing identifiable people. Thus, we retained 248 images from KonIQ-10k and 537 from Pixabay. All images have a minimum width of $2048$ pixels and are annotated at their original resolution.
@@
In this way, we account for the non-linearity of the slider scale. The final value $\Omega(I)$ represents the ground-truth IIS value of the image $I$, or MOIS.
S2.4 Effects of Variable Viewing Conditions
Viewing conditions – such as display characteristics (contrast, brightness), environmental lighting, image presentation (cropping), and the viewer’s visual acuity or viewing distance – significantly influence perceived image quality at an individual level. Laboratory-based experiments offer tight control over these factors but often result in atypical and low-variability conditions that may not generalize to real-world scenarios. In contrast, crowdsourcing introduces diverse and less controlled viewing conditions, which can lead to lower participant reliability. However, when reliable group opinions are obtained through crowdsourcing, these assessments are more likely to generalize to a broader population under similar conditions. Therefore, despite variability in display conditions, the group-wise average opinion for IIS effectively captures perceptual IIS measurements. As detailed in Secs. 4.2 and S2.5, our annotation setup is highly reliable. Consequently, our crowdsourcing approach not only maintains high reliability but also ensures that the collected data is representative of real-world viewing conditions, thereby enhancing the validity and applicability of our findings.
S2.5 Reliability of the IIS Annotations
In Sec. 4.2 we demonstrated the reliability of the annotations in our dataset by comparing the inter-group agreement for the IISA-DB and popular NR-IQA datasets. In this section, we further assess the reliability of our ground-truth annotations by computing the confidence intervals (CIs) for the aggregated IIS of each image (MOIS). Following a well-established methodology, we sample 20 opinions with replacement for each image, compute their geometric mean, and repeat this process 100 times. The 95% CI is then estimated as half the range between the $2.5^{\text{th}}$ and $97.5^{\text{th}}$ percentiles of the simulated geometric means. This yields an average CI of $0.057$ for our dataset. As a reference, we consider the KonX NR-IQA dataset, whose annotation conditions were similar to those of our dataset, involving remote studies and freelance participants. The average CI of the MOS for the KonX dataset is $0.046$. Given the differences between the IISA and NR-IQA tasks, some variation in the CIs is expected. Therefore, this outcome indicates that our annotation process has a level of reliability comparable to a highly reliable NR-IQA dataset such as KonX.
S2.6 Examples of Image-IIS Pairs
@@
Following the evaluation protocol described in Sec. 6.1, we evaluate the performance of this zero-shot multi-scale approach on the proposed IISA-DB dataset. We employ two versions of the TOPIQ [5] model pre-trained on the SPAQ [9] and UHD-IQA [15] datasets, which feature high-resolution images similar to those in IISA-DB. We use $n_{s}=100$ scales and report the results in Tab. S1. We observe that the zero-shot multi-scale approach achieves unsatisfactory performance, regardless of the pre-training dataset. We attribute this outcome to the fact that NR-IQA models struggle to handle the change in perceptual quality caused by downscaling, as noted in [34, 16]. In addition, the multi-scale approach requires multiple model forward passes to obtain a single IIS prediction, which can be inefficient.

### v2

@@ Acknowledgements
In Sec. 4.2, we discuss the precision of subjective measurements, indicated by the confidence intervals of the aggregated annotations. The precision levels for IISA and NR-IQA are comparable. However, the leverage factor $\gamma>1$ implies that minor changes in quality result in larger variations in scale. Therefore, achieving a specific precision in measuring scale equates to an even finer precision in measuring quality. Thus, $\gamma$ acts as a leverage or magnifying factor. This indicates that the sensitivity of our annotation tool for IIS is higher when detecting subtle differences in quality compared to traditional quality rating scales, making IISA suitable for fine-grained quality assessment.
@@
S1.4 IISA Applications: Additional Details
We present scenarios across industry and research where IISA is the perfect tool to optimize quality and resolution trade-offs.
Printing and publishing Printing an image too large can accentuate flaws present in the source image while printing too small sacrifices detail. IISA can guide the selection of print dimensions and resolution (dot-per-inch, DPI). This ensures consistently high-quality prints. Traditionally, this task is managed by an expert operator. IISA automates these decisions, allowing scalable deployment in online printing systems and enabling non-expert users to make optimal choices independently.
Moreover, web developers and UI designers often need to serve images across devices with different screen sizes and resolutions. Typically, responsive design uses fixed rules, whereas IISA enables content-awareness. For instance, an online image gallery can automatically size each photo based on its intrinsic scale. This ensures that users see images at the best quality for their device while saving bandwidth and load time.
Gaming and graphics rendering Modern games employ dynamic resolution scaling to maintain high FPS. However, choosing the amount of rescaling on each axis can dramatically affect quality. IISA offers a principled solution to this problem. Moreover, if the target FPS is not fixed, a game engine could automatically downscale rendered frames (slightly degraded by aliasing or motion blur) until just before quality starts dropping, ensuring players get the clearest visuals with optimal performance (FPS).
Computational photography Smartphone cameras rely on computational photography to balance resolution and noise. Although sensors may reach 100+MP, phones often merge pixels in low light to produce cleaner lower MP images – effectively searching for the image’s intrinsic scale. IISA makes this process explicit and optimal. In tasks like super-resolution or denoising, algorithms can use IISA to determine when further resolution or noise reduction stops improving quality.
Benchmarking IQA methods Traditional IQA metrics predict quality at a fixed resolution, while IISA requires accuracy at multiple scales. Our experiments show that off-the-shelf IQA models – NR-IQA trained on traditional datasets – perform poorly on the IISA task. By directly evaluating alignment with human perception across scales, IISA serves as a valuable benchmark. Performance improvements of NR-IQA methods on the IISA-DB benchmark indicate a deeper understanding of the quality–resolution trade-off, which is critical for both academic research and real-world image processing.
Extending IQA study methodology IISA introduces new methods for subjective image quality evaluation. Traditional IQA often has viewers rate an image’s quality at a fixed resolution, which can be difficult for subtle degradations in no-reference settings. By contrast, IISA asks viewers to resize an image until it “looks best”, inherently comparing quality across scales. This approach improves sensitivity to minor artifacts (see Sec. S1.1).
Training dataset curation for image restoration Constructing image datasets requires managing images of varying quality. Traditionally, low-quality images are discarded to avoid introducing erroneous priors into restoration models. However, IISA provides a more nuanced approach: rather than discarding low-quality images, it downscales them to their intrinsic scale to maximize quality. This preserves content diversity because discarding images purely on quality can disproportionately exclude dynamic or low-light scenes, which tend to be blurrier and noisier. By applying IISA, we mitigate that bias while maintaining a wider range of content.
@@
Figure S4: Distribution of the width of the images of the IISA-DB dataset at their original and intrinsic scale.
@@
Aiming for higher quality in the Pixabay subset, we selected newer camera models from a list of 71, focusing on those released after 2010 with full-frame sensors. From a pool of over one million images on Pixabay with EXIF information, we filtered for photos with a width greater than 4,000 pixels, resulting in approximately 18,000 images that met all criteria. We sampled for diversity 600 images, maintaining uniform distributions among binned normalized favorites, likes, downloads, and user tags using a method similar to [33]. Most of these were captured with Canon EOS 1/5/6D Mark 2/3/4 cameras, using various lenses and capture settings. For the KonIQ-10k subset, we also sampled for diversity regarding quality levels and machine tags – with a confidence greater than 80% – using the same stratified procedure.
The last filtering step we applied was removing images containing identifiable people. Thus, we retained 248 images from KonIQ-10k and 537 from Pixabay. All images have a minimum width of $2048$ pixels and are annotated at their original resolution. Fig. S4 illustrates the distribution of image widths at their original and intrinsic scales.
@@
In this way, we account for the non-linearity of the slider scale. The final value $\Omega(I)$ represents the ground-truth IIS value of the image $I$, or MOIS. Across the 785 images composing our dataset, the average MOIS is 0.347, with per-image MOIS ranging from 0.060 to 0.811.
S2.4 Examples of Image-IIS Pairs
@@
Following the evaluation protocol described in Sec. 6.1, we assess the performance of this zero-shot multi-scale approach on the proposed IISA-DB dataset. We employ two versions of the TOPIQ [5] model pre-trained on the SPAQ [9] and UHD-IQA [15] datasets, which feature high-resolution images similar to those in IISA-DB. We use $n_{s}=100$ scales and report the results in Tab. S1. We observe that the zero-shot multi-scale approach achieves unsatisfactory performance, regardless of the pre-training dataset. We attribute this to NR-IQA models struggling to handle the change in perceptual quality caused by downscaling, as noted in [33, huang2024high]. In addition, the multi-scale approach requires multiple model forward passes to obtain a single IIS prediction, which can be inefficient.
@@
S5 Future Work
Our work suggests several promising directions for future research, including: 1) extending our dataset to incorporate more types of images, such as super-resolved, synthetically distorted, and computer-generated images; 2) conducting additional subjective studies – similar to those of KonX – to futher validate our assumption related to how image quality changes with scale; 3) analyzing the impact of different types of distortion on the IIS.